OpenStudy (anonymous):

The length of a rectangle is 1m more than twice its width, and the area of the rectangle is 28m^2 . Find the dimensions of the rectangle.

OpenStudy (jdoe0001):

so the Area of a rectangle is Length x Width so the Length is whatever the Width is +1 Width = w Length = w+1 thus the Area of Length x Width => \(\bf w \times (w+1) = 28\) solve for "w" to get the Width to get the Length, just add 1 to that

OpenStudy (jdoe0001):

hh, darn I hate that, I just notice futher material

OpenStudy (jdoe0001):

o the Area of a rectangle is Length x Width so the Length is whatever the Width TWICE that + 1 Width = w Length = 2w+1 thus the Area of Length x Width => \(\bf w×(2w+1)=28\) solve for "w" to get the Width to get the Length, just add 1 to that